OpenStudy (anonymous):

y varies directly as x. If y = 12 when x = 8, find y when x = 6

OpenStudy (anonymous):

y=kx Find "k" 12=k*8 12/8=k k=1.5 Equation y=1.5x Find x when y=6 6=1.5x 6/1.5=x x=4

OpenStudy (anonymous):

thanks (:

OpenStudy (anonymous):

Your welcome :)